# Spectrality of Prime Size Tiles

### v5

@@ 3 Main results
Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be the linear transform defined as
$Ve_{i}=\begin{cases}v_{i},&\text{if }i=1,\ldots,p-1;\\ e_{i},&\text{if }i>p-1.\end{cases}$
$V$ has full rank since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions. Further set
@@
By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$. This implies that $A$ is tiling in $\mathbb{Z}^{d}$, and its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$. ∎

### v6

@@ 3 Main results
Without loss of generality we may assume that one of these points is the origin, and we denote the other p1p-1 points by $v_{1},\ldots,v_{p-1}$. Let $V$ be an arbitrary linear transform that maps $e_{i}$ to $v_{i}$ for i<=p1i<=p-1, such $V$ exists since $0,v_{1},\ldots,v_{p-1}$ are in general linear positions (which means that $v_{1},\ldots,v_{p-1}$ are linearly independent). Further set
@@
By Corollary 1, $E$ is tiling in $\mathbb{Z}^{p-1}$, thus $A=VE$ is also tiling in the space obtained by applying $V$ to $\operatorname{span}(E)$, which is a sublattice since $\operatorname{span}(E)=\mathbb{Z}^{p-1}$. This implies that $A$ is also tiling in $\mathbb{Z}^{d}$. Its tiling complement can be produced explicitly: If $E^{\prime}$ is a tiling complement of $E$ in $\mathbb{Z}^{p-1}$, and $F$ is the set formed by taking one representative from each coset of $V\mathbb{Z}^{p-1}$ (viewed as an additive subgroup) in $\mathbb{Z}^{d}$, then $B=(VE^{\prime})\oplus F$ is a tiling complement of $A$. ∎